OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (zzr0ck3r):

can you solve 7a^2-14a+2=-5?

OpenStudy (anonymous):

isn't it 7a^2 -14a +7 = 0 now?

OpenStudy (anonymous):

Hi Death :) thanks for viewing my question

OpenStudy (anonymous):

well im here to help now. sorry for leaving earlier. had to finish some summer school homework

OpenStudy (anonymous):

No problem! Thank you so much!

OpenStudy (anonymous):

I hate my Pre-Calc course...the text book is the freaking online textbook is the most difficult thing to navigate ever...plus it's a college course while I'm in high school math! I have no choice in the matter...damn pre-calc...

OpenStudy (anonymous):

its either ''a = -0.286 - 0.958'' or ''a = -0.286 + 0.958".

OpenStudy (anonymous):

I feel you. I hate online class textbooks

OpenStudy (anonymous):

so by either do you mean both or you're not sure? I don't THINK I have to use unit circle for this one...probably...

OpenStudy (anonymous):

because it doesn't give me an interval

OpenStudy (anonymous):

wait? no interval?

OpenStudy (zzr0ck3r):

7a^2 -14a +7 = 0 yes. solve this 7(a-1)^2 = 0 so a = 1 so sin(x) =1 this is at pi/2,6pi/2,10pi/2 = pi/2+2pi*k where k is an integer

OpenStudy (zzr0ck3r):

there are infinite solutions, no interval needed.

OpenStudy (anonymous):

no, no interval...just all possible solutions

OpenStudy (zzr0ck3r):

7a^2 -14a +7 = 0 yes. solve this 7(a-1)^2 = 0 so a = 1 so sin(x) =1 this is at pi/2,5pi/2,9pi/2 so solutions are pi/2+2pi*k where k is an integer

OpenStudy (anonymous):

where did k come from? Ummm I got confused...

OpenStudy (anonymous):

he means the "k" should be the integer

OpenStudy (zzr0ck3r):

well sin(x) = 1 at pi/2 and also (since we are not restricted to an interval) at 5pi/2 and 9pi/2 so {....,pi/2,5pi/2,9pi/2,......} another way to write this is pi/2+2pi*k where k is an integer

OpenStudy (zzr0ck3r):

for k=0 you get pi/2+2pi*0 = pi/2 for k = 1 you get pi/2+2pi*1 = 5pi/2 for k = 2 you get 9pi/2 .......

OpenStudy (zzr0ck3r):

so if you let k run through all the integers {...-3,-2,-1,0,1,2,3.....} you will get "all" of the soltions

OpenStudy (zzr0ck3r):

I know this is allot, but I think you should spend some time and make sure you understand everything I said. It will help you allot in this class..

OpenStudy (anonymous):

okay thank you very much for all your time